# A Systematic Literature Review of Parameter-Efficient Fine-Tuning for Large Code Models

### v2

@@ Abstract.
The rise of Artificial Intelligence (AI)-and particularly Large Language Models (LLMs) for code–has reshaped Software Engineering (SE) by enabling the automation of tasks such as code generation, bug detection, and repair. However, these models require significant computational resources for training and fine-tuning, posing challenges for real-world adoption in resource-constrained environments. To address this, the research community has increasingly turned to Parameter-Efficient Fine-Tuning (PEFT)–a class of techniques that enables the adaptation of large models by updating only a small subset of parameters, rather than the entire model. In this Systematic Literature Review (SLR), we examine the growing application of PEFT techniques–across a wide range of software engineering tasks. We analyze how these methods are used to optimize various deep learning (DL) architectures, focusing on their impact on both performance and efficiency. Our study synthesizes findings from 27 peer-reviewed papers, identifying patterns in configuration strategies and adaptation trade-offs. The outcome of this review is a comprehensive taxonomy that categorizes PEFT usage by task type, distinguishing between generative (e.g., Code Summarization) and non-generative (e.g., Code Clone Detection) scenarios. Our findings aim to inform future research and guide the practical deployment of PEFT in sustainable, AI-powered software development.
@@ 2.1. Large Code Models
The effectiveness of these models is further augmented by continuous improvements in both architecture and task-specific tuning methods. Innovations such as infilling objectives (Li et al., 2023a) and instruction tuning (Xu et al., 2024) have been integrated into models like StarCoder and CodeLlama, improving their performance on tasks such as Code Completion and Code Editing (Roziere et al., 2023). Additionally, LCMs have been effectively applied to domain-specific and low-resource programming languages, including Verilog (Liu et al., 2023a) and Haskell (van Dam et al., 2024), often utilizing transfer learning from high-resource languages like Python and Java to overcome data scarcity issues (Zhang et al., 2023b). Beyond their conventional applications in programming tasks, LCMs have also been leveraged in emerging areas such as Autonomous Agent Systems for workflow automation (Hong et al., 2024) and in assessing the quality of AI-generated code (Wei et al., 2023a).
@@ 3. Methodology: Study Selection and Analysis
  • RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods? This research question aims to identify the types of base models and architectural families that are commonly fine-tuned using PEFT techniques. Additionally, it explores how different PEFT methods are instantiated across various model types (e.g., encoder-only, decoder-only, encoder-decoder) to support software engineering tasks. The goal is to understand which architectural configurations are most “compatible” with PEFT optimization methods.
@@
      Snowballing and Manual Addition of Studies: Identified and included two additional studies through backward snowballing, resulting in a final set of 27 papers.
@@ 3.2.2. Snowballing and Manual Addition of Studies
After incorporating these two additional studies, our final study set comprised 27 papers. The distribution of these papers across publication venues is shown in Fig. 2.
@@ 3.3. Data Extraction and Synthesis
Having finalized our corpus of 27 primary studies through rigorous search, filtering, and snowballing procedures (as described in Section 3.1 and Section 3.2), we proceeded to extract and organize data to answer our research questions. In the following subsections, we first describe our data extraction process and then outline the synthesis approach, including how we grouped software engineering tasks into generative and non-generative categories.
@@ 3.3.2. Data Extraction
To systematically address our three research questions, we followed a structured extraction protocol that enabled systematic collection of relevant attributes from each of the 27 selected primary studies.
@@
[…]
Overall, our review encompasses 27 studies covering 17 distinct SE tasks, which we classify into eleven (11) generative and six (6) non-generative task groups. These are further detailed in Table 1, where we report the task category in the leftmost column, followed by a full description of the software engineering task type–adopting the terminology established by Watson et al. (Watson et al., 2022). Finally, the third column reports the corresponding studies in which each task type was identified.
@@
Code Generation (CG) Generates source code from natural language descriptions or formal specifications. (Wang et al., 2025; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Choi and Lee, 2023; Chen et al., 2023b; Shi et al., 2023; Ayupov and Chirkova, 2022; Yang et al., 2024a)
@@
Code Translation (CTrans) Converts code from one programming language to another while preserving functionality. (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a)
@@
Code Completion (CComp) Predicts and completes the next code token or line based on the surrounding context. (Shi et al., 2023; Liu et al., 2024a)
@@ 4.1. RQ1: Which software engineering tasks have been explored using PEFT methods?
This research question investigates and quantifies how PEFT methods are applied to enhance the efficiency, effectiveness, or both, of software engineering tasks. From our analysis of 27 primary studies, we identified 17 distinct SE tasks where PEFT-based techniques have been applied. These tasks have already been categorized into generative and non-generative types in Section 3.3.1. Generative tasks typically involve creating or transforming code artifacts (e.g., Code Generation, Code Summarization, or Automated Program Repair), while non-generative tasks focus on retrieval, detection, or classification of code properties (e.g., Code Search, Defect Detection). We further organize these 17 SE tasks based on their input-output modalities as follows: NL2Code tasks—which involve generating code from natural language, such as Code Generation and Unit Test Generation—appear in 10 studies; Code2NL tasks—which involve translating code into natural language, such as Code Summarization, Commit Message Generation, and Code Review Generation—appear in 15 studies; and Code2Code tasks—which involve code-to-code transformations or analysis, such as Code Translation, Code Refinement, Automated Program Repair, Code Clone Detection, and others—are explored in 21 studies.
We obtained 16 papers where more than one SE task was supported via PEFT methods (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Chen et al., 2023b; Saberi and Fard, 2023; Fan et al., 2024; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Goel et al., 2022; Liu et al., 2024a).
@@
This approach allows us to build a taxonomy that accurately reflects the original goal and scope of each PEFT application. Between 2022 and 2025, research on PEFT in software engineering has grown significantly, with an increasing number of papers addressing diverse tasks. This growing trend is depicted in Fig. 4 and highlights an early traction of Code Summarization (13 studies), and Code Clone Detection (6 studies), while Code Generation (9 studies) has only seen a notable rise in PEFT supports recently. Nonetheless, since our collection of relevant studies concluded on February 2025, we expect that additional papers-particularly in the areas of software engineering and broader applications-will be published by the end of the current year (2025).
Upon closer examination of generative tasks, we observe that Code Summarization accounted for 44.2% (13 studies) and Code Generation with 9 studies (33.3%) are the most frequently explored tasks, reflecting strong interest in improving model efficiency for synthesis-based application. Prefix Tuning and Prompt Tuning are among the earliest and most extensively tested methods for adapting Transformer-based models to Code Summarization with minimal trainable parameters. Adapter-based techniques are also popular, as evidenced in (Chen et al., 2023b; Ayupov and Chirkova, 2022; Liu et al., 2023b; Wang et al., 2023a).
@@
Moving forward, Code Translation also shows growing traction with 7 different papers exploring the extent to which PEFT techniques, including Additive, Reparametrized, as well as Hybrid, can serve the purpose of effective and efficient fine-tuning. For example, Liu et al. (Liu et al., 2023b) evaluates multiple PEFT methods including Adapter, LoRA, Prefix Tuning, Prompt Tuning, and Mix-And-Match (MAM) on CodeT5 and PLBART to translate code segments from Java to C#. In a different but related study, Wang et al. (Wang et al., 2023c) explore the capabilities of Prompt Tuning with varying sizes of CodeT5.
@@
Other explored tasks include Just-in-time Comment Update, Method Name Recommendation, Code Completion, and Unit Test Generation. Across these tasks, LoRA, Prefix Tuning, and Adapters remain the most frequently used methods, often applied to large-scale models such as CodeLlama (Roziere et al., 2023), CodeT5 (Wang et al., 2021), and InCoder (Fried et al., 2022).
@@
Moving beyond generative tasks, PEFT techniques have also been explored for non-generative SE applications, with Code Clone Detection being the most prominent among them, appearing in 6 studies (22.2%). To this end, a relevant approach is Telly-K, proposed by Shi et al. (Shi et al., 2023), which selectively tunes the bottom layers of UniXcoder and GraphCodeBERT to detect Java code clones on the BigCloneBench dataset (Svajlenko and Roy, 2016) which is Defect Detection is another important non-generative task where PEFT methods have shown promise, appearing in 5 studies. This task focuses on identifying buggy code, often using C or Java datasets like Devign (Zhou et al., 2019) and JIT-Defects4J. For example, Liu et al. (Liu et al., 2024b) explores Adapter Tuning and LoRA on a diverse set of models—including CodeBERT, GraphCodeBERT, and PLBART using the JIT-Defects4J dataset with over 16K training instances to detect Just-in-time Software Defects. Similarly, Wang et al. (Wang et al., 2023c) applies Prompt Tuning to CodeBERT and CodeT5 variants on the Devign dataset, which includes over 27K C code snippets with nearly half labeled as defective. These efforts highlight how PEFT enables effective adaptation for real-world bug detection scenarios with limited parameter updates.
@@
PEFT techniques have been applied to 17 diverse SE tasks, with a strong focus on generative tasks and growing interest in non-generative ones. While tasks like Code Summarization and Generation remain the most studied, several emerging tasks have also seen initial exploration. Common PEFT methods include Prompt Tuning, Adapter, LoRA, (IA)3, and QLoRA, applied across popular models such as CodeT5, CodeLlama, and CodeBERT. These findings demonstrate the adaptability and effectiveness of PEFT in software engineering with different model architectures and varying sizes.
@@ 4.2. RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods?
Before detailing the results of this analysis, we remind the reader that one paper can employ more than one techniques, as a results of that, we found that encoder-decoder architectures are the most prevalent in PEFT-based software engineering optimization, appearing in 29 software engineering task instances across 12 studies (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Xie et al., 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Peng et al., 2024; Zlotchevski et al., 2022). This is followed by encoder-only architectures with 22 applications across 10 studies (Hajipour et al., 2024; Saberi and Fard, 2023; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c, 2022b, a; Goel et al., 2022), while we found that auto-regressive models (i.e., decoder-only architectures) are employed to automate 18 software engineering activities across 11 studies (Li et al., 2024b; Liu et al., 2024a; Fakih et al., 2024; Yang et al., 2024a; Hajipour et al., 2024; Fan et al., 2024; Yu et al., 2024; Liu et al., 2024c; Afrin et al., 2025; Huang et al., 2025; Wang et al., 2025). These findings indicate a clear preference for encoder-decoder architectures among researchers, likely due to their effectiveness in handling both understanding and generation tasks common in SE applications.
@@
When examining specific model variants, the CodeLlama family shows particularly flexible adoption, with CodeLlama-7B appearing in 7 SE instances across 5 studies (Li et al., 2024b; Fakih et al., 2024; Fan et al., 2024; Liu et al., 2024c; Afrin et al., 2025), CodeLlama-13B in 4 SE instances across 2 studies (Li et al., 2024b; Fan et al., 2024), and CodeLlama-34B in 4 studies (Liu et al., 2024a; Fakih et al., 2024; Afrin et al., 2025; Huang et al., 2025). Crossing tasks and models, we find out that CodeLlama-7B is used across diverse tasks including Automated Program Repair, Code Generation, Commit Message Generation, and Code Summarization. CodeLlama-13B extends this usage to Code Completion, Code Translation, and Unit Test Generation, while CodeLlama-34B supports more resource intensive tasks such as Code Completion and Automated Program Repair. This shows that the CodeLlama models are effective across a wide range of tasks, including NL2Code tasks like Code Generation, Code2NL tasks like Code Summarization and Code Review Generation.
@@
Several recent studies have explored a range of model architectures in their evaluations, including StarCoder-15B (Huang et al., 2025; Liu et al., 2024a), DeepSeek-Coder variants (Li et al., 2024b; Liu et al., 2024a; Afrin et al., 2025), and QWen-14B (Liu et al., 2024a), reflecting interest in evaluating both general purpose language models and those specialized for code related tasks. The architecture distribution reveals clear task-specific patterns. Encoder-decoder models (CodeT5, PLBART) maintain their dominance in generation-heavy tasks and thus they are naturally applied in the context of generative applications. On the other hand, encoder-only models (CodeBERT, GraphCodeBERT) remain preferred for analysis-oriented tasks. Finally, decoder-only models (particularly CodeLlama variants) are demonstrating growing adoption across both Code Generation and Automated Program Repair tasks, as evidenced by multiple studies (Li et al., 2024b; Liu et al., 2024c; Wang et al., 2025; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Fan et al., 2024). Decoder-only models appeared exclusively in generative tasks. Specifically, they are used in NL2Code tasks such as Code Generation (4 studies), Code2NL tasks such as Code Summarization, Commit Message Generation, and Code Review Generation (each reported in 1 study), and Code2Code tasks such as Automated Program Repair (2 studies), Code Translation, Code Refinement, and Just-in-Time Comment Update (each reported in 1 study). The evolution of model selection shows a marked progression from earlier focus on established architectures to current exploration of larger, more specialized models. This shift is particularly evident in the rising adoption of 13B+ parameter models like CodeLlama variants (7B/13B/34B), DeepSeek-Coder-33 (Liu et al., 2024a; Afrin et al., 2025), and LLaMA-2-70B (Liu et al., 2024a), enabled by advancing PEFT techniques that make such models more accessible even in resource-constrained environments.

### v3

@@ Abstract.
The rise of Artificial Intelligence (AI)-and particularly Large Language Models (LLMs) for code–has reshaped Software Engineering (SE) by enabling the automation of tasks such as code generation, bug detection, and repair. However, these models require significant computational resources for training and fine-tuning, posing challenges for real-world adoption in resource-constrained environments. To address this, the research community has increasingly turned to Parameter-Efficient Fine-Tuning (PEFT)–a class of techniques that enables the adaptation of large models by updating only a small subset of parameters, rather than the entire model. In this Systematic Literature Review (SLR), we examine the growing application of PEFT techniques–across a wide range of software engineering tasks. We analyze how these methods are used to optimize various deep learning (DL) architectures, focusing on their impact on both performance and efficiency. Our study synthesizes findings from 28 peer-reviewed papers, identifying patterns in configuration strategies and adaptation trade-offs. The outcome of this review is a comprehensive taxonomy that categorizes PEFT usage by task type, distinguishing between generative (e.g., Code Summarization) and non-generative (e.g., Code Clone Detection) scenarios. Our findings aim to inform future research and guide the practical deployment of PEFT in sustainable, AI-powered software development.
@@
These authors contributed equally to this work.
@@ 2.1. Large Code Models
The effectiveness of these models is further augmented by continuous improvements in both architecture and task-specific tuning methods. Innovations such as infilling objectives (Li et al., 2023a) and instruction tuning (Xu et al., 2024) have been integrated into models like StarCoder and CodeLlama, improving their performance on tasks such as Code Completion and Code Editing (Roziere et al., 2023). Additionally, LCMs have been effectively applied to domain-specific and low-resource programming languages, including Verilog (Liu et al., 2023a) and Haskell (van Dam et al., 2024), often utilizing transfer learning from high-resource languages like Python and Java to overcome data scarcity issues (Zhang et al., 2023c). Beyond their conventional applications in programming tasks, LCMs have also been leveraged in emerging areas such as Autonomous Agent Systems for workflow automation (Hong et al., 2024) and in assessing the quality of AI-generated code (Wei et al., 2023a).
@@ 3. Methodology: Study Selection and Analysis
  • RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods?
    This research question aims to identify the types of base models and architectural families that are commonly fine-tuned using PEFT techniques. Additionally, it explores how different PEFT methods are instantiated across various model types (e.g., encoder-only, decoder-only, encoder-decoder) to support software engineering tasks. The goal is to understand which architectural configurations are most “compatible” with PEFT optimization methods.
@@
      Snowballing and Manual Addition of Studies: Identified and included two additional studies through backward snowballing, resulting in a final set of 28 papers.
@@ 3.2.2. Snowballing and Manual Addition of Studies
After incorporating these two additional studies, our final study set comprised 28 papers. The distribution of these papers across publication venues is shown in Fig. 2.
@@ 3.3. Data Extraction and Synthesis
Having finalized our corpus of 28 primary studies through rigorous search, filtering, and snowballing procedures (as described in Section 3.1 and Section 3.2), we proceeded to extract and organize data to answer our research questions. In the following subsections, we first describe our data extraction process and then outline the synthesis approach, including how we grouped software engineering tasks into generative and non-generative categories.
@@ 3.3.2. Data Extraction
To systematically address our three research questions, we followed a structured extraction protocol that enabled systematic collection of relevant attributes from each of the 28 selected primary studies.
@@
RQ3 (Performance and Efficiency): We focused on comparing performance and computational efficiency between PEFT-optimized large code models and fully fine-tuned baseline models. For each study, we identified the best-performing PEFT methods (multiple methods if present) and recorded whether reported outcomes indicated improvements ( P E), equivalence ( P E), or declines ( P E) in performance and efficiency compared to the full fine-tuning baseline. Efficiency metrics such as reduced training time and memory consumption were noted where available. These findings were synthesized into a categorical summary, enabling task-wise comparisons of PEFT impacts on performance and efficiency.
Overall, our review encompasses 28 studies covering 19 distinct SE tasks, which we classify into twelve (12) generative and seven (7) non-generative task groups. These are further detailed in Table 1, where we report the task category in the leftmost column, followed by a full description of the software engineering task type–adopting the terminology established by Watson et al. (Watson et al., 2022). Finally, the third column reports the corresponding studies in which each task type was identified.
@@
Code Generation (CG) Generates source code from natural language descriptions or formal specifications. (Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Choi and Lee, 2023; Chen et al., 2023b; Shi et al., 2023; Ayupov and Chirkova, 2022; Yang et al., 2024a), (Wang et al., 2025a)
@@
Code Translation (CTrans) Converts code from one programming language to another while preserving functionality. (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a), (Wang et al., 2025a)
@@
Code Completion (CComp) Predicts and completes the next code token or line based on the surrounding context. (Shi et al., 2023; Liu et al., 2024a), (Wang et al., 2025a)
@@
Protocol Buffer Transformation (PBT) Generates code functions that transform data structures between different format specifications. (Wang et al., 2025a)
@@
Header File Prediction (HFP) Predicts the correct file paths for header dependencies given source file information. (Wang et al., 2025a)
@@ 4.1. RQ1: Which software engineering tasks have been explored using PEFT methods?
This research question investigates and quantifies how PEFT methods are applied to enhance the efficiency, effectiveness, or both, of software engineering tasks. From our analysis of 28 primary studies, we identified 19 distinct SE tasks where PEFT-based techniques have been applied. These tasks have already been categorized into generative and non-generative types in Section 3.3.1. Generative tasks typically involve creating or manipulating code-related artifacts as output (e.g., source code, code documentation), often in an open-ended or token-by-token fashion. This includes tasks like code summarization, code generation, and commit message generation, even if the latter involves some form of templating–because the model must still synthesize content beyond merely selecting from fixed classes. On the other hand, non-generative tasks are those where the output space is bounded or categorical, such as classification, detection, or retrieval. For example, code classification–even though it produces an output–falls under the non-generative category because it selects from a finite set of labels and does not require constructing novel sequences of tokens. We further organize the 12 generative tasks based on their input-output modalities as follows: NL2Code tasks–which involve generating code from natural language, such as Code Generation and Unit Test Generation–appear in 12 studies; Code2NL tasks–which involve translating code into natural language, such as Code Summarization, Commit Message Generation, and Code Review Generation–appear in 16 studies; and Code2Code tasks–which involve code-to-code transformations or analysis, such as Code Translation, Code Refinement, Automated Program Repair, and others–are explored in 21 studies. While generative tasks are categorized based on their input-output modalities due to the diversity and structural implications of the output (i.e., text or code), non-generative tasks do not require such fine-grained modality breakdown. This is because the output across non-generative tasks is typically constrained to label spaces (i.e., binary, multi-class, regression values). To enhance clarity and better reflect the structural diversity of non-generative tasks, we classify them according to standard machine learning output paradigms (Zhang et al., 2023b; Peña and Herbold, 2025). Specifically, we group them into the categories as follows – Classification: Tasks where the model predicts a discrete class label, such as Code Clone Detection, Defect Detection and Code Classification; Regression: Tasks involving continuous-valued outputs; Ranking: Tasks that return ordered lists or scores, such as Code Search or Traceability Link Recovery.
We obtained 17 papers where more than one SE task was supported via PEFT methods (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Chen et al., 2023b; Saberi and Fard, 2023; Fan et al., 2024; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Goel et al., 2022; Liu et al., 2024a; Wang et al., 2025a)..
@@
This approach allows us to build a taxonomy that accurately reflects the original goal and scope of each PEFT application. Between 2022 and 2025, research on PEFT in software engineering has grown significantly, with an increasing number of papers addressing diverse tasks. This growing trend is depicted in Fig. 4 and highlights an early traction of Code Summarization (13 studies), and Code Clone Detection (6 studies), while Code Generation (10 studies) has only seen a notable rise in PEFT supports recently. Nonetheless, since our collection of relevant studies concluded on February 2025, we expect that additional papers-particularly in the areas of software engineering and broader applications-will be published by the end of the current year (2025).
Upon closer examination of generative tasks, we observe that Code Summarization accounted for $\sim$46.4% (13 studies) and Code Generation with 10 studies ($\sim$35.7%) are the most frequently explored tasks, reflecting strong interest in improving model efficiency for synthesis-based application. Prefix Tuning and Prompt Tuning are among the earliest and most extensively tested methods for adapting Transformer-based models to Code Summarization with minimal trainable parameters. Adapter-based techniques are also popular, as evidenced in (Chen et al., 2023b; Ayupov and Chirkova, 2022; Liu et al., 2023b; Wang et al., 2023a).
@@
Moving forward, Code Translation also shows growing traction with 8 different papers exploring the extent to which PEFT techniques, including Additive, Reparametrized, as well as Hybrid, can serve the purpose of effective and efficient fine-tuning. For example, Liu et al. (Liu et al., 2023b) evaluates multiple PEFT methods including Adapter, LoRA, Prefix Tuning, Prompt Tuning, and Mix-And-Match (MAM) on CodeT5 and PLBART to translate code segments from Java to C#. In a different but related study, Wang et al. (Wang et al., 2023c) explore the capabilities of Prompt Tuning with varying sizes of CodeT5.
@@
Other explored tasks include Just-in-time Comment Update, Method Name Recommendation, Code Completion, Protocol Buffer Transformation and Unit Test Generation. Across these tasks, LoRA, Prefix Tuning, and Adapters remain the most frequently used methods, often applied to large-scale models such as CodeLlama (Roziere et al., 2023), CodeT5 (Wang et al., 2021), and InCoder (Fried et al., 2022).
@@
Moving beyond generative tasks, PEFT techniques have also been explored for non-generative SE applications, with Code Clone Detection being the most prominent among them, appearing in 6 studies ($\sim$21.4%). Analyzing the distribution of non-generative tasks across machine learning paradigms reveals clear patterns. Classification tasks dominate, with 16 studies covering Cloze Test (2), Code Clone Detection (6), Code Review (1), Defect Detection (5), Method Name Consistency Check (1), and Header File Prediction (1). Ranking tasks appear in only 3 studies, all in Code Search applications, and no regression-based tasks were identified. Overall, approximately $\sim$84% of non-generative PEFT applications focus on discrete classification problems, split into 8 papers on binary classification ($\sim$66.7%), 2 on multi-class classification ($\sim$16.7%), and 2 addressing both ($\sim$16.7%). To this end, a relevant approach is Telly-K, proposed by Shi et al. (Shi et al., 2023), which selectively tunes the bottom layers of UniXcoder and GraphCodeBERT to detect Java code clones on the BigCloneBench dataset (Svajlenko and Roy, 2016) which is Defect Detection is another important non-generative task where PEFT methods have shown promise, appearing in 5 studies. This task focuses on identifying buggy code, often using C or Java datasets like Devign (Zhou et al., 2019) and JIT-Defects4J. For example, Liu et al. (Liu et al., 2024b) explores Adapter Tuning and LoRA on a diverse set of models—including CodeBERT, GraphCodeBERT, and PLBART using the JIT-Defects4J dataset with over 16K training instances to detect Just-in-time Software Defects. Similarly, Wang et al. (Wang et al., 2023c) applies Prompt Tuning to CodeBERT and CodeT5 variants on the Devign dataset, which includes over 27K C code snippets with nearly half labeled as defective. These efforts highlight how PEFT enables effective adaptation for real-world bug detection scenarios with limited parameter updates.
@@
PEFT techniques have been applied to 19 diverse SE tasks, with a strong focus on generative tasks and growing interest in non-generative ones. While tasks like Code Summarization and Generation remain the most studied, several emerging tasks have also seen initial exploration. Common PEFT methods include Prompt Tuning, Adapter, LoRA, (IA)3, and QLoRA, applied across popular models such as CodeT5, CodeLlama, and CodeBERT. These findings demonstrate the adaptability and effectiveness of PEFT in software engineering with different model architectures and varying sizes. Task concentration patterns reveal that Code Summarization represents $\sim$46.4% of studies while Code Generation accounts for $\sim$35.7%, indicating clear research priorities in synthesis-based applications. Cross-task versatility analysis shows that $\sim$60.7% of studies apply PEFT to multiple SE tasks, demonstrating broader methodological flexibility than single-task approaches. The distribution across task categories shows non-generative tasks like Code Clone Detection representing $\sim$21.4% of studies, revealing systematic coverage of both generative and non-generative software engineering activities.
@@ 4.2. RQ2: Which deep learning models and architectures are commonly optimized with PEFT methods?
Before detailing the results of this analysis, we remind the reader that one paper can employ more than one techniques, as a results of that, we found that encoder-decoder architectures are the most prevalent in PEFT-based software engineering optimization, appearing in 29 software engineering task instances across 12 studies (Hajipour et al., 2024; Choi and Lee, 2023; Zhu et al., 2023; Xie et al., 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b, 2023a; Peng et al., 2024; Zlotchevski et al., 2022). This is followed by encoder-only architectures with 22 applications across 10 studies (Hajipour et al., 2024; Saberi and Fard, 2023; Saberi et al., 2024; Shi et al., 2023; Liu et al., 2024b, 2023b; Wang et al., 2023c, 2022b, a; Goel et al., 2022), while we found that auto-regressive models (i.e., decoder-only architectures) are employed to automate 19 software engineering activities across 12 studies (Li et al., 2024b; Liu et al., 2024a; Fakih et al., 2024; Yang et al., 2024a; Hajipour et al., 2024; Fan et al., 2024; Yu et al., 2024; Liu et al., 2024c; Afrin et al., 2025; Huang et al., 2025; Wang et al., 2025b, a). These findings indicate a clear preference for encoder-decoder architectures among researchers, likely due to their effectiveness in handling both understanding and generation tasks common in SE applications. It is important to note that our goal in examining model architectures is not to suggest that architecture type prescriptively determines the choice of PEFT method. Instead, this analysis aims to document how current practices reveal consistent patterns in how PEFT techniques are applied across model families. By empirically mapping the distribution of PEFT methods over different architectures, we capture how practical considerations–such as model structure, parameter scale, and task orientation–correlate with the adoption of particular strategies.
@@
When examining specific model variants, the CodeLlama family shows particularly flexible adoption, with CodeLlama-7B appearing in 7 SE instances across 5 studies (Li et al., 2024b; Fakih et al., 2024; Fan et al., 2024; Liu et al., 2024c; Afrin et al., 2025), CodeLlama-13B in 6 SE instances across 3 studies (Li et al., 2024b; Fan et al., 2024; Wang et al., 2025a), and CodeLlama-34B in 5 studies (Liu et al., 2024a; Fakih et al., 2024; Afrin et al., 2025; Huang et al., 2025; Wang et al., 2025a). Crossing tasks and models, we find out that CodeLlama-7B is used across diverse tasks, including Automated Program Repair, Code Generation, Commit Message Generation, and Code Summarization. CodeLlama-13B extends this usage to Code Generation, Code Completion, Code Translation, and Unit Test Generation, Protocol Buffer Transformation, while CodeLlama-34B supports more resource intensive tasks such as Code Completion and Automated Program Repair. This shows that the CodeLlama models are effective across a wide range of tasks, including NL2Code tasks like Code Generation, Code2NL tasks like Code Summarization and Code Review Generation.
@@
Several recent studies have explored a range of model architectures in their evaluations, including StarCoder-15B (Huang et al., 2025; Liu et al., 2024a), DeepSeek-Coder variants (Li et al., 2024b; Liu et al., 2024a; Afrin et al., 2025; Wang et al., 2025a), and QWen-14B (Liu et al., 2024a), reflecting interest in evaluating both general purpose language models and those specialized for code related tasks. The architecture distribution reveals clear task-specific patterns. Encoder-decoder models (CodeT5, PLBART) maintain their dominance in generation-heavy tasks, and thus, they are naturally applied in the context of generative applications. On the other hand, encoder-only models (CodeBERT, GraphCodeBERT) remain preferred for analysis-oriented tasks. Finally, decoder-only models (particularly CodeLlama variants) are demonstrating growing adoption across both Code Generation and Automated Program Repair tasks, as evidenced by multiple studies (Li et al., 2024b; Liu et al., 2024c; Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Fan et al., 2024). Decoder-only models appeared exclusively in generative tasks. Specifically, they are used in NL2Code tasks such as Code Generation (5 studies), Code2NL tasks such as Code Summarization, Commit Message Generation, and Code Review Generation (each reported in 1 study), and Code2Code tasks such as Automated Program Repair (2 studies), Code Translation, Code Refinement, Protocol Buffer Transformation and Just-in-Time Comment Update (each reported in 1 study). The evolution of model selection shows a marked progression from earlier focus on established architectures to current exploration of larger, more specialized models. This shift is particularly evident in the rising adoption of 13B+ parameter models like CodeLlama variants (7B/13B/34B), DeepSeek-Coder-33 (Liu et al., 2024a; Afrin et al., 2025; Wang et al., 2025a), and LLaMA-2-70B (Liu et al., 2024a), enabled by advancing PEFT techniques that make such models more accessible even in resource-constrained environments.